OpenStudy (anonymous):

Simplify: 1.6 X 10^3 ____________ 4 X 10^-2

OpenStudy (anonymous):

1600/0.04 = 40,000 = 40X10^3

OpenStudy (anonymous):

It's all one equation, it's a fraction I just didn't know how to type it

OpenStudy (anonymous):

it doesnt have any variables... you can simplify it down to a real number. 40,000

OpenStudy (anonymous):

thank you I get it now!

OpenStudy (anonymous):

also how do you solve: -3(a^2b^-1)^3 {-------------}^-2 7(a^-5b^6) The brackets should cover both equations.

OpenStudy (anonymous):

note that when you take the exponent of an exponent you add ex: (2^2)^2 = 2^4 right off i see the a^2^3 up there which turns into a^5th and you have an a^-5 down in the numerator. start there, and see if that helps any

OpenStudy (anonymous):

okay thank you so much!